OpenStudy (anonymous):

Which shows ^6√24^2 in simplified radical form with the smallest possible index?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi c:

OpenStudy (misty1212):

start by dividing by two and get \[\sqrt[3]{24}\]

OpenStudy (misty1212):

then since \(24=8\times 3\) you get \[\sqrt[3]{24}=\sqrt[3]{3}\times \sqrt[3]{3}=2\sqrt[3]{3}\]

OpenStudy (misty1212):

oops i meant \[\sqrt[3]{24}=\sqrt[3]{8}\times \sqrt[3]{3}=2\sqrt[3]{3}\]

OpenStudy (anonymous):

Oh, okay. I understand now. Thank you for the help. ^-^

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]